Math 291 Lab 8 Due Monday March 19, 2018

Instructions: Turn in your lab by emailing it to fagerstrom@mnstate.edu. You should email both your raw TEX (.tex) file and your compiled document (in .pdf format).

You will be graded on both your raw TEX code and the accuracy of your compiled document. Name your file in the format of JaneDoeLab8.tex

• Create a document that has: – The ‘report’ document class

– At least two chapters

– At least two sections in each chapter

– At least three subsections within at least one of the sections

– A couple of subsubsections (they may or may not be in the same subsection, it is up to you)

– A couple of paragraphs (they may or may not be in the same region of your document, it is up to you)

– Have a title (which includes the phrase Math 291 L^{A}TEX Lab 8), author (you), and date.

– Have a title page, a table of contents, and a list of figures.

– Uses each of the pictures from my website that were not used in either the presentation or the Example document from class. They are:

∗ CatsLoveReading

∗ CatTree

∗ Kittens ∗ MarbleInez ∗ PuzzleCat ∗ safety ∗ SunnyCatPile ∗ SunnyFlower ∗ Warmth

Note: You can reuse any of the other pictures if you want as well, or include any of your own. But you must use these nine pictures.

– Resizes at least one picture using all three of the width, height, and keepaspectratio commands.

– Resizes at least one picture using the width and height commands in such a way that it is clearly either squashed or elongated.

– Rotates some of the pictures ∗ Rotates CatTree so that it is correctly rotated.

∗ Rotates another pictures so that it is at a 60 degree angle.

– Has at least five figures that have numbers and captions.

∗ At least one of the figure should include multiple graphics in a single figure, like my figure that had three pictures of Marble in it in the example file.

∗ At least one should be (intentionally) at the top of a page.

∗ At least one should be ‘here’ (or at least intended to be ‘here’).

– Has at least one unnumbered figure (use the ⋆ version of the figure environment). This figure should still have a caption.

– Has language that refers to several different chapters, sections, subsections, or figures.

∗ There should be at least six labels and references, with at least one of each of the four types.

∗ Add the appropriate verbiage to your reference, so that you are referencing a Chapter or Figure that has a certain number, not just the number. [See page 24 of the 27 in the presentation for ideas if you need them.]

– The document should not just be a bunch of figures — add some appropriate paragraphs (even if they are short paragraphs), as in the example document.